# Generalized Linear Bandits with Limited Adaptivity

### v3

@@ A.2 Optimal Design Guarantees
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2d^{2}.$
@@
$\mathbb{P}\left\{3\varepsilon n{\mathbf{I}}+\sum_{i=1}^{n}x_{i}x_{i}^{\mathsf{T}}\succeq\frac{n}{8}\ \operatorname*{\mathbb{E}}_{x\sim{\mathcal{D}}}\left[xx^{\mathsf{T}}\right]\right\}\geq 1-2d\exp\left(\frac{-\varepsilon n}{8}\right)$
@@
${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{G}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
@@
We note that the analysis of Ruan et al. [2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}}^{2}$. We obtain such a bound here and use it in the analysis.
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{k}}\right)\quad\forall k\in[M]$ (17)
@@
$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\geq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
We obtain the first to inequalities, (16) and (17), by a direct use of Corollary A.16. For (18) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M). Since, $T\geq d^{32}\log{2T}^{2}$ the event stated in Lemma A.13 holds with probability greater than $1-\frac{1}{T^{2}}$. ∎
@@ Appendix B Regret Analysis of RS-GLinCB
Recall ${\mathcal{T}}_{o}$ denotes the set of warm-up rounds. We write $\tau_{w}$ to denote the size of the set $\tau_{w}=|{\mathcal{T}}_{o}|$. We define the following (scaled) data matrix
@@
At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the warm-up rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
$\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
@@
$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{w}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
@@
Next, we obtain a bound on the number of warm-up rounds.
@@
Algorithm 2, during its entire execution, enters the enters the warmup segment at most $2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$ times.
Proof.
Via the warm-up condition (Line 6) $\lVert x\rVert_{\mathbf{V}^{-1}}^{2}>1/(R^{2}\kappa\gamma^{2})$. Let $\mathbf{V}_{m}$ be the sequence of $\mathbf{V}$ matrices (line 7 of 2) for $m\in{\mathcal{T}}_{o}$. That is, $\mathbf{V}_{1}=\lambda{\mathbf{I}},\mathbf{V}_{m}=\sum_{s\in[m-1]\setminus{\mathcal{T}}_{o}}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$). Thus,
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\geq\frac{\tau_{w}}{R^{2}\kappa\gamma^{2}}$ (19)
Furthermore, by the Elliptic Potential Lemma (Lemma B.11) we have
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{\tau_{w}}{\lambda d}\right)$ (20)
Combining (20) and (19) we have
$\tau_{w}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{w}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
We define ${\mathcal{E}}_{w}$ be the event define in Lemma B.2, that is, ${\mathcal{E}}_{w}=\{\lVert\widehat{\theta}_{w}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma_{t}\}$.
Lemma B.4.
If in round $t$ the warmup criteria is not satisfied and event ${\mathcal{E}}_{w}$, holds, we have
@@
Now we define the following quantities:
@@
$\Theta$ =$=\left\{\theta|\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
@@
Let $\delta\in(0,1)$. Then, under event ${\mathcal{E}}_{w}$, with probability $1-\delta$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta\coloneqq\textbf{c}\sqrt{d\log(T/\delta)}$.
@@
We have for all main rounds $s\in[\tau-1]\setminus{\mathcal{T}}_{o}$,
@@
If round $t$ is a main round for Algorithm 2, then, under event ${\mathcal{E}}_{\tau}$, $\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}$
@@
For main rounds in Algorithm 2, let $x_{t}$ be the played arm defined via line 14 Algorithm 2. Then, under event ${\mathcal{E}}_{\tau}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}~{}.$
@@
For non-warm up rounds, the arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$ (b) for $x_{t},x^{*}_{t}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
@@
Suppose $x^{\prime}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}LCB_{w}(x)$. Now, we have, for all $x\in{\mathcal{X}}_{t}$,
@@
Thus, for $x^{*}_{t}$, we can write,
$\langle x^{*}_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{*}_{t},\theta^{*}\rangle$
$\geq\langle x^{\prime},\theta^{*}\rangle$ (optimality of $x^{*}_{t}$)
$\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
Thus, $x^{*}_{t}\in{\mathcal{X}}^{\prime}_{t}$.
Since $x_{t}$ is also in ${\mathcal{X}}^{\prime}_{t}$,
$\langle x_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
$\geq\langle x_{t}^{*},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ ($x^{\prime}$ has max LCB)
@@
$\implies\langle x^{*}_{t}-x_{t},\theta^{*}\rangle$ $\leq 2\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}+2\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$
Finally, since this is a non warm-up round, $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}<\frac{1}{R\gamma\sqrt{\kappa}}$ for all $x\in{\mathcal{X}}_{t}$. Plugging this above,
@@
B.1 Proof of Theorem 4.2
We prove Theoreom 4.2 in this section
@@
$\leq R\tau_{w}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for warm-up rounds)
@@
$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Elliptic Potential Lemma; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
@@
B.2 Proof of Lemma 4.1
We first restate Lemma 4.1.
@@
Note that in Algorithm 2, policy changes happen only in the warm-up rounds and when the determinant of ${\mathbf{H}}_{t}$ doubles in the main round. Thus, total number of policy switches is upper bounded by the number of warm-up rounds and the number of times $\det({\mathbf{H}}_{t})$ doubles. The first quantity is bounded by Lemma B.3 while the second quantity is bounded by Lemma B.14. Thus in total, the number of policy changes in Algorithm A is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
B.3 Some Useful Lemmas

### v4

@@ A.5 Optimal Design Guarantees
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}}^{2}\right]\leq 2\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}}\left[\max_{x\in{\mathcal{X}}}\ \left\lVert x\right\rVert_{{\mathbf{W}}^{-1}_{G}}^{2}\right]\leq 2d^{2}.$
@@
$\mathbb{P}\left\{3\varepsilon n{\mathbf{I}}+\sum_{i=1}^{n}x_{i}x_{i}^{\mathsf{T}}\succeq\frac{n}{8}\ \operatorname*{\mathbb{E}}_{x\sim{\mathcal{D}}}\left[xx^{\mathsf{T}}\right]\right\}\geq 1-2d\exp\left(-\frac{\varepsilon n}{8}\right)$
@@
${\mathbf{H}}_{k}\succeq\frac{\tau_{k}}{8}\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\operatorname*{\mathbb{E}}_{\widetilde{x}\sim\pi_{k}({\mathcal{X}})}\widetilde{x}\widetilde{x}^{\mathsf{T}}.$
@@
We note that the analysis of Ruan et al. [2021] gives an optimal guarantee (in expectation) on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}$, but not on $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}^{2}$. We obtain such a bound here and use it in the analysis.
@@
$\operatorname*{\mathbb{E}}_{{\mathcal{X}}\sim{\mathcal{D}}_{k}}\max_{x\in{\mathcal{X}}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}^{2}$ $\leq O\left(\frac{d^{2}}{\tau_{k}}\right)\quad\forall k\in[M]$ (19)
@@
$\left\lVert\widetilde{x}\right\rVert_{{\mathbf{H}}_{k}^{-1}}\leq\frac{8}{\tau_{1}}\left\lVert\widetilde{x}\right\rVert_{{\mathbf{W}}_{k}^{-1}}$
We obtain the first two inequalities, (18) and (19), by a direct use of Corollary A.16. For (20) we note that for every phase we have at least $O(\sqrt{T})$ samples for learning the Distributional optimal design policy (for any M). Since, $T\geq d^{32}\log{2T}^{2}$ the event stated in Lemma A.13 holds with probability greater than $1-\frac{1}{T^{2}}$. ∎
@@ Appendix B Regret Analysis of RS-GLinCB
Recall ${\mathcal{T}}_{o}$ denotes the set of rounds when switching criterion I is satisfied. We write $\tau_{o}$ to denote the size of the set $\tau_{o}=|{\mathcal{T}}_{o}|$. We define the following (scaled) data matrix
@@ B.1 Confidence Sets for Switching Criterion I rounds
At any round $t$, let $\widehat{\theta}_{o}$ be the maximum likelihood estimate calculated using set of rewards observed in the rounds ${\mathcal{T}}_{o}$. With probability at least $1-\delta$ we have
$\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma.$
@@
$+(1+2RS)\left(\frac{\sqrt{R^{2}\lambda}}{2}+\frac{2d}{\sqrt{R^{2}\lambda}}\log\left(1+\frac{\tau_{o}}{R^{2}\lambda d}\right)+\frac{2}{\sqrt{R^{2}\lambda}}\log\left(\frac{1}{\delta}\right)+\frac{2}{\sqrt{R^{2}\lambda}}d\log(2)\right)$
@@
B.2 Confidence Sets for non-Switching Criterion I rounds
We define ${\mathcal{E}}_{w}$ be the event defined in Lemma B.2, that is, ${\mathcal{E}}_{w}=\{\lVert\widehat{\theta}_{o}-\theta^{*}\rVert_{{\mathbf{H}}^{*}_{w}}\leq\gamma\}$.
@@ B.2 Confidence Sets for non-Switching Criterion I rounds
If in round $t$ the switching criteria I is not satisfied and the event ${\mathcal{E}}_{w}$ holds, we have
@@
Recall that $\tau$ is the round when the Switching Criterion II (Line 9) is satisfied. Now we define the following quantities:
@@
$\Theta$ =$=\left\{\theta:\left\lVert\theta-\widehat{\theta}_{o}\right\rVert_{\mathbf{V}}\leq\gamma\sqrt{\kappa}\right\}$
@@
Let $\delta\in(0,1)$. Then, under event ${\mathcal{E}}_{w}$, with probability $1-\delta$, $\left\lVert\widehat{\theta}_{\tau}-\theta^{*}\right\rVert_{{\mathbf{H}}^{*}_{\tau}}\leq\beta$.
@@
We have for all rounds $s\in[\tau-1]\setminus{\mathcal{T}}_{o}$,
@@
B.3 Bounding the instantaneous regret
In this subsection, we will only consider rounds $t\in[T]$ which does not satisfy Switching Criterion I. Let $x_{t}\in{\mathcal{X}}_{t}$ be the played arm defined via line 14 Algorithm 2. Further, let $x^{*}_{t}\in{\mathcal{X}}_{t}$ be the best available arm in that round.
@@ B.3 Bounding the instantaneous regret
Under the event ${\mathcal{E}}_{\tau}$, for all $x\in{\mathcal{X}}_{t}$, we have, $\lvert\langle x,\widehat{\theta}_{\tau}-\theta^{*}\rangle\rvert\leq\beta\lVert x\rVert_{{\mathbf{H}}_{\tau}^{-1}}~{}.$
@@
Under event ${\mathcal{E}}_{\tau}$, $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq 2\sqrt{2}\beta\left\lVert x_{t}\right\rVert_{{\mathbf{H}}^{*-1}_{t}}~{}.$
@@
The arm set ${\mathcal{X}}^{\prime}_{t}$ obtained after eliminating arms from ${\mathcal{X}}_{t}$ (line 13 Algorithm 2), under event ${\mathcal{E}}_{w}$, satisfies: (a) $x^{*}_{t}\in{\mathcal{X}}_{t}$, (b) $\langle x^{*}_{t}-x_{t},\theta^{*}\rangle\leq\frac{4}{R}$
@@
Suppose $x^{\prime}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}LCB_{o}(x)$. Now, we have, for all $x\in{\mathcal{X}}_{t}$,
@@
Thus, $UCB_{o}(x^{*}_{t})=\langle x^{*}_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}\geq\langle x^{*}_{t},\theta^{*}\rangle\geq\langle x^{\prime},\theta^{*}\rangle\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}=LCB_{o}(x^{\prime})~{},$ where the second inequality is due to optimality of $x^{*}_{t}$. Hence, $x^{*}_{t}$ is not eliminated, implying $x^{*}_{t}\in{\mathcal{X}}^{\prime}_{t}$. This completes the proof of (a).
Since $x_{t}$ is also in ${\mathcal{X}}^{\prime}_{t}$ (by definition),
$UCB_{o}(x_{t})=\langle x_{t},\widehat{\theta}_{o}\rangle+\gamma\sqrt{\kappa}\left\lVert x_{t}\right\rVert_{\mathbf{V}^{-1}}$ $\geq\langle x^{\prime},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{\prime}\right\rVert_{\mathbf{V}^{-1}}$
$\geq\langle x_{t}^{*},\widehat{\theta}_{o}\rangle-\gamma\sqrt{\kappa}\left\lVert x^{*}_{t}\right\rVert_{\mathbf{V}^{-1}}$ ($x^{\prime}$ has max $LCB_{o}(\cdot)$)
@@
[…]
Finally, since Switching Criterion I is not satisfied in this round, $\left\lVert x\right\rVert_{\mathbf{V}^{-1}}<\frac{1}{R\gamma\sqrt{\kappa}}$ for all $x\in{\mathcal{X}}_{t}$. Plugging this above,
@@
B.4 Proof of Theorem 4.2
In this subsection, we complete the proof of the regret bound of RS-GLinCB(Algorithm 2). We first restate Theoreom 4.2 and then prove it. For every round $t\in[T]$, we use $x_{t}\in{\mathcal{X}}_{t}$ to denote the arm played by the algorithm and $x^{*}_{t}$ to denote the best available arm in that round.
Theorem B.10 (Theorem 4.2).
@@
$\leq R\tau_{o}+\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\mu\left(\langle x^{*}_{t},\theta^{*}\rangle\right)-\mu\left(\langle x_{t},\theta^{*}\rangle\right)$ (Upper bound of $R$ for rounds in ${\mathcal{T}}_{o}$)
@@
$\leq 2e^{5}\sqrt{2}\beta\sqrt{\left(\sum_{t\in[T]\setminus{\mathcal{T}}_{o}}\dot{\mu}\left(\langle x^{*}_{t},\theta^{*}\rangle\right)\right)\cdot 2d\log\left(1+\frac{RT}{\lambda d}\right)}$ (Lemma B.12; $\left\lVert\tilde{x}_{t}\right\rVert_{2}\leq R$)
@@
B.5 Bounding number of policy updates: Proof of Lemma 4.1
We first obtain a bound on the number of rounds when Switching Criterion I is satisfied. Then we restate Lemma 4.1 and present its proof. Here, we use ${\mathcal{T}}_{o}$ to denote the collection of all rounds till $T$ for which Switching Criterion I is satisfied.
Lemma B.11.
Algorithm 2, during its entire execution, satisfies the Switching Criterion I at most $2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$ times.
Proof.
Recall that Switching Criterion I (Line 4) is satisfied, when $\lVert x\rVert_{\mathbf{V}^{-1}}^{2}>1/(R^{2}\kappa\gamma^{2})$ for some $x\in{\mathcal{X}}_{t}$. Let $\mathbf{V}_{m}$ be the sequence of $\mathbf{V}$ matrices (line 6 of Algorithm 2) for $m\in{\mathcal{T}}_{o}$. That is, $\mathbf{V}_{1}=\lambda{\mathbf{I}},\mathbf{V}_{m}=\sum_{s\in[m-1]\cap{\mathcal{T}}_{o}}x_{s}x_{s}^{\intercal}+\lambda{\mathbf{I}}$. In these rounds, by Line 5 of Algorithm 2, we have that the arm played $x_{t}$ is such that $x_{t}=\operatorname*{arg\,max}_{x\in{\mathcal{X}}_{t}}\left\lVert x\right\rVert_{\mathbf{V}_{t}^{-1}}$. Therefore,
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x_{t}\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\geq\frac{\tau_{o}}{R^{2}\kappa\gamma^{2}}$ (22)
Furthermore, by the Elliptic Potential Lemma (Lemma B.12) we have
$\sum_{t\in{\mathcal{T}}_{o}}\left\lVert x_{t}\right\rVert_{\mathbf{V}_{t}^{-1}}^{2}\leq 2d\log\left(1+\frac{\tau_{o}}{\lambda d}\right)$ (23)
Combining (23) and (22) we have
$\tau_{o}\leq 2dR^{2}\kappa\gamma^{2}\log{\left(1+\frac{\tau_{o}}{\lambda d}\right)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T)}\leq 2dR^{2}\kappa\gamma^{2}\log{(T/\delta)}$
@@
Note that in Algorithm 2, policy changes happen only in the rounds when either of the Switching Criteria are triggered. The number of times Switching Criterion I is triggered is bounded by Lemma B.11. On the other hand, the number of times Switching Criterion II is triggered is equal to the number of times determinant of ${\mathbf{H}}_{t}$ doubles, which is bounded by Lemma B.15. Thus in total, the number of policy changes in Algorithm 2 is upper bounded by $2dR^{2}\kappa\gamma^{2}\log(T/\delta)+\textbf{c}d\log(T)~{}.$
B.6 Some Useful Lemmas